evaluate ∫ sin^2(x) / 1 + cos x dx

evaluate
∫  sin^2(x) / 1 + cos x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question

I = \int \frac{sin^{2}x}{1+cosx}dx
I = \int \frac{1-cos^{2}x}{1+cosx}dx
I = \int \frac{(1-cosx)(1+cosx)}{1+cosx}dx
I = \int (1-cosx)dx
I = \int dx-\int cosxdx
I = x - sinx + c